### v3

# Digital currency hardware wallets and the essence of money

Abstract
Many proposals for the design and implementation of digital wallets assume that the purpose of the wallet is to enable offline payments via custodial accounts, ignoring the real problems faced by individuals and businesses that engage in retail payments, such as the anticompetitive behaviour of payment platforms and the decline of cash. More importantly, the proposals ignore the raison d’être of digital currency as a kind of digital money that can be held independently of custodians. Finally, the proposals demonstrate a profound lack of imagination about the nature of digital money and the devices that could be used to hold, manage, and exchange it. From these presumptions flows a set of architectural requirements that stifle the promise of digital currency to deliver novel and efficient ways to exchange value in the digital economy. In this article, we critically assess the essential problems that digital currency solutions are being proposed to solve, particularly with respect to the future of payments and the future of cash. We assess the validity of common justifications for account-based payments and certified hardware in the context of alternative designs, limitations, and trade-offs. We conclude that the interests of consumers would be better served by design approaches to digital currency that anticipate that digital assets would be held outside accounts, stored offline, but transacted online, without requiring the use of trusted hardware.
1 Introduction
The proliferation of cashless payments has given rise to a growing preponderance of merchants who refuse to accept cash, as well as the suggestion that private-sector payment systems could subvert the authority of central banks to implement monetary policy. This development has, in turn, focussed the attention of central banks and other financial institutions around the world on digital currency, particularly the idea that central banks could issue central bank digital currency (CBDC) for retail use. In principle, retail CBDC would be a direct obligation of the central bank, like cash or central bank reserves, but existing in digital form and held by ordinary consumers and non-financial businesses.
Many proponents of retail CBDC have argued that it must be “cash-like” to deliver value to end-users. This argument is at least partially right. For example, it might be expected that a cash-like system would afford users privacy by design, resistance to discrimination and profiling, and the ability to possess and control the assets that they use to engage with the economy. These are certainly desirable properties. However, some authorities have interpreted the argument that a digital payment system must be cash-like to mean that a digital payment system must enable dual-offline transactions, namely those without connectivity of any kind between either of the counterparties and any third parties, in contrast to “single-offline” transactions, wherein a mutually trusted third party is reachable by one of the counterparties. We suggest that this argument constitutes a leap of logic that must be examined carefully, starting with its implicit assumptions that any new payment system must be appropriate for all scenarios, that existing solutions such as cash are insufficient to handle payment scenarios in which neither party has access to a network connection, and that there is a specific need for future digital currency infrastructure to enable a kind of digital transaction that is not commonly available to users today.
Importantly, the idea that digital currency transactions must mirror the current system for electronic retail payments demonstrates an assumption that the main promise of digital currency concerns accessing custodial accounts whilst offline. But this assumption ignores decades of history of digital currency, wherein the most groundbreaking systems, including not only Bitcoin but also DigiCash two decades earlier, drew attention by introducing a way to hold money outside of custodial accounts. The idea that central banks would never support such an architecture ignores the fact that central banks have always supported the issuance and circulation of tokens outside accounts, in the form of cash. Within this framework, offering cash in digital form is the next logical step.
Dual-offline transactions impose additional security requirements to prevent double-spending attacks, in addition to the new security requirements applicable to digital bearer instruments in general. Some prominent experts have proposed “trusted computing” [1, 2] as a solution to the security problem, reasoning that if the security of bearer instruments is a cause for concern, then perhaps there is benefit to be realised in making an effort to ensure that the devices that can hold them must meet a certain standard. After all, many users are accustomed to carrying certified hardware, and the prevalence of certified hardware among common user devices might be seen as a justification for their use in the context of digital currency. For example, smart cards commonly rely upon tamper-resistant chips containing secrets issued by their manufacturers [3]. Mobile phones often contain secure enclaves to restrict what their users can do with their devices, for example to prevent users from copying multimedia files.
We also consider the conflation of technology, which can be freely and directly deployed by an individual or business to perform a function, with services, which are functions performed by third parties. In a world of smartphone applications for which most of the computation is nearly always performed by remote service providers, it is easy to ignore this distinction. However, since the ability to possess and control money directly is essential to its function, we must be careful to distinguish between technology that is under the control of the end-user and technology that is under the control of a third party.
Of course, the fact that certain methods are commonly used in some contexts does not justify their use in others, and it certainly does not justify mandating the use of such methods in the context of future infrastructure that intrinsically must be available for general use by the public. It is all too easy for those in the payments industry to assume that everyone has or should have a smartphone and a bank account. Mandating the use of certified hardware carries many costs, and the benefits of certified hardware are not as clear as its proponents claim.
The aforementioned considerations raise the question of how end-users will hold digital currency assets, and in particular, what the role of a wallet should be. ISO 22739:2024 defines a wallet as an “application or mechanism used to generate, manage, store or use private keys and public keys or other digital assets” [42]. The simplicity of this definition would seem to understate the potentially complex features required to implement token management and custody, an exploration of which is given by Barresi and Zatti [5]. We address some of the assumptions that Barresi and Zatti make about the design of wallets and suggest different ways of viewing the problem. In particular, we critically address the arguments that trusted computing is an indispensable enabler of digital currency solutions, that digital currency must be held in accounts, that wallets must be linked to the identity of a user, and that a wallet is properly seen as “not a place where you can store cash [but] that helps you keep track of it” [5].
The rest of the article is organised as follows. In the next section, we consider the arguments in favour of offline payments, toward an understanding of precisely what is meant by the use of the term “offline”, as well as the specific kinds of offline payments for which trusted computing has been proposed as a necessary solution and the underlying motivations for this reasoning. In Section 3, we critically examine the requirement for trusted computing in the context of central bank digital currency, identifying several key problems that undermine its perceived value. In Section 4, we argue that token-based architectures are appropriate for privacy and self-custody. In Section 5, we discuss the role of the issuer as distinct from the trusted third parties that facilitate transactions. In Section 6, we offer an alternative approach that does not rely upon trusted computing and argue that this and similar approaches are better-suited for retail digital currency infrastructure in the general case. In Section 7, we summarise our arguments.
2 The sophistry of dual-offline payments
Recently, several prominent organisations, including the International Monetary Fund [6] as well as central banks such as the Bank of England [7] and the Bank of Canada [8], have released reports suggesting that dual-offline payments are desirable or even essential to the functionality of a future retail CBDC system. Such bold statements are striking, considering the prevailing lack of consensus about what the requirements for digital currency should be in general. In this section, we shall critically explore the argument that the purpose of digital currency is to facilitate access to custodial accounts in a dual-offline payment scenario.
2.1 Kinds of payments
First, we must ask: What exactly constitutes an “offline” payment? We consider the most common kinds of retail payment mechanisms and assess them in terms of whether they are online or offline:
  • E-commerce transactions. Transactions that involve consumers using personal computers or mobile devices to interact with websites require the internet by definition. They represent a significant and growing share of the total value consumers spend in the economy, and naturally, they require consumers to be online.
  • Bank transfers. Bank transfers are mechanisms for transferring money from a sender’s account with a sending bank to a recipient’s account with a receiving bank. As account-holders, both the sender and the recipient must be identified. Bank transfers can be arranged online, or an account-holder can arrange a bank transfer via a conversation with a bank teller or a computer controlled by the bank. The sending bank debits the sender’s account, and the receiving bank credits the recipient’s account. Settlement between the banks is managed via an interbank clearing and settlement network. In most cases, interbank settlement networks cannot guarantee “immediate” settlement within a few seconds, and bank transfers are usually not considered appropriate at the point of sale or for transactions that require acceptance in real-time.
  • Electronic card payments at the point of sale (POS). Electronic card payments provide a mechanism by which merchants can accept in-person payments from consumers in real-time. The consumer’s bank (the issuer) and the merchant’s bank (the acquirer) both participate in a payment network. The issuer provides a card to the consumer, and the acquirer offers to connect a merchant’s POS device to the payment network. The card serves as a form of identification: Its purpose is to specify a particular bank account and to identify its user as the beneficiary of that account. To protect against unauthorised copying of the identification credential, cards are usually designed with features that provide tamper-resistance. The client is not assumed to have an Internet connection, and the POS device provides a way for the card to authenticate itself to the payment network. If the authentication is successful, then the payment network effectively underwrites the transaction, implicitly providing credit to the issuer and prompting the acquirer to accept the transaction immediately. The payment network informs the merchant that the transaction is approved, and the merchant can then accept the payment in real-time. Merchants pay interchange fees to support this mechanism.
  • Closed-loop payment systems at the point of sale. With closed-loop payment networks, both consumers and merchants are primary customers of a platform operator, which maintains accounts with balances for all of its customers. When a consumer pays a merchant in-person, the platform operator debits the consumer’s account and credits the merchant’s account. For a transaction to succeed, either the consumer or the merchant must have an active connection to the platform operator. The consumer must authenticate successfully to the platform operator, although this can be done via the merchant’s device, as it is with card payment networks. Crucially, banks are not generally involved in transactions; the platform operator manages the process.
  • Cash. Cash comprises a set of physical tokens, each of which is a direct obligation of a central bank or monetary authority. Cash is held directly by individuals and non-financial businesses as a store of value, and it is also held by banks to service withdrawals, wherein banks debit accounts of customers in exchange for providing them with cash. Cash is fungible, which is to say that individual cash tokens are mutually substitutable and undifferentiated in practice. Since the consummation of a cash transaction is the exchange of a physical token, cash transactions between parties are offline by definition. Regulators generally do not have a direct view into cash transactions, and assessment of regulatory and tax compliance depends upon accounting and reporting.
E-commerce transactions are certainly online, and cash transactions are certainly offline. However, whether point of sale transactions and bank transfers are online or offline is a matter of perspective. For example:
  • Is it enough for one party but not both to be connected to the Internet?
  • Is it necessary to be able to reach a particular service run by a specific operator, or is it sufficient to be able to reach any one of a set of operators?
  • Can a transaction to be provisionally allocated using the network but consummated when both parties are disconnected from the network?
Notwithstanding some measure of understandable confusion and inconsistency in the use of the term “offline” accompanying calls for considering offline payments to be an essential feature of digital currency, it can be said that some measure of online technology supports virtually all retail payments today, excepting only payments in cash, which has seen a dramatic decline in use throughout the world over the past decade.
With this in mind, it is reasonable to ask: What problem do proponents of retail CBDC intend to solve, for which dual-offline payments constitute a better solution than online payments involving assets held offline? We acknowledge that there are some limited real-world use cases for dual-offline payments, such as payments aboard transportation systems that lack reliable network connectivity throughout the system [9]. However, such examples are rare indeed, owing to the prevalence of network connectivity in locations suitable for digital transactions. Furthermore, we note that payment scenarios similar to this example do not actually require dual-offline payments; their requirements could also be satisfied by time-shifting, in which consumers purchase e-money tokens offline and consummate the transactions offline by selectively revealing proofs of purchase [10]. And so, it is fair to ask: Might the idea of dual-offline payments constitute a solution in search of a problem, an ephithet that a report by the UK House of Lords prominently ascribed to CBDC in general [11]?
2.2 Digital currency use cases
One might argue that conducting transactions using digital currency held by custodians is, from the perspective of users and counterparties, not much different than conducting transactions practice of using ordinary bank accounts, with a possible exception of a different regulatory environment for the digital currency custodian. The custodian is still a gatekeeper that could fail to meet its obligations; the user is still bound to the custodian by identification requirements and still does not hold the assets directly. One might also argue that digital currency held by custodians is even less useful than bank deposits, if it is assumed that digital currency tokens cannot be rehypothecated and therefore cannot contribute to money creation.
We might reasonably conclude that non-custodial wallets, which allow users to hold digital assets directly, are a foundational benefit of digital currency. However, even if we accept that non-custodial wallets are essential to digital currency, the idea that offline payments are necessary for digital currency is certainly false. Users have an interest to possess and control digital assets directly, regardless of the context in which they spend them. Holding digital assets directly takes custodians out of outbound transactions, allowing users to avoid the risk that the custodians might extract fees, block certain (or all) transactions from taking place, or build profiles of the behaviour of their customers by monitoring the details of their transactions, such as time, location, and counterparty. Bearer wallets can be used in a variety of contexts, including but not limited to in-person POS transactions. Importantly, the benefits of non-custodial wallets can also be realised in a fully online context such as e-commerce.
Proponents of digital currency are right to acknowledge the secular decline in the use of cash in retail consumer transactions as a primary motivator for digital currency. As fewer consumers use cash for transactions, the variable revenues associated with operating a cash infrastructure are falling below the fixed costs. This pattern is evident in much of the developed world. For example, in the UK, banks are closing ATMs and branches [12], making cash more difficult for consumers to withdraw [13]. Some merchants have decided that the marginal cost of the various expenses related to accepting cash payments, such as tills, employee time, cash accounts with banks, and the means of transporting cash, outweighs the benefit of providing goods and services to cash-only customers. As a result, the value of cash to consumers is reduced, since customers no longer have assurance that their cash will be accepted at the point of sale, creating a feedback loop.
And so, when proponents of digital currency say that digital currency should allow “cash-like” transactions for retail users of money, they are referencing the idea that digital currency can provide individuals with a payment option that provides the key affordances of cash, including direct ownership, privacy, and equal access without discrimination. Users of digital currency would hold direct obligations of the central bank, and, given the right regulations, they might have reason to believe that their digital currency would be accepted by retailers that require electronic transactions.
Digital currency can certainly address problems associated with the systematic decline in the set of available options for consumers to access cash and the increasing tendency for merchants to refuse payments in cash at the point of sale. Users can load non-custodial wallets with the help of a personal device with network access, as well as via facilities provided by banks or ATM networks. Consumers can use non-custodial wallets in-person, to interface with point of sale devices in the same manner that they use debit cards or other custodial payment mechanisms today. Such payments can be considered “offline” both in the sense that they are in-person and in the sense that that the non-custodial wallets themselves do not require Internet connections. We refer to such payment scenarios as single-offline, because only one of the counterparties to a transaction has access to third parties via a network or other means of connectivity. Single-offline payments include in-person EMV card payments via POS systems and are therefore constitute a significant share of the payments that have replaced cash payments in the retail payments sector over the past twenty years.
2.3 Hold offline, transact online
The assumption that transactions that are not between accounts must be offline is wrong. This assumption would imply that all payments that are not online between payment accounts must be offline between peers that are disconnected from real-time communication with any other party. In particular, the proposal by the European Central Bank (ECB) for the digital euro [14] holds that these two models for payment, online via accounts and dual-offline between peers, are comprehensively exhaustive, but this is simply false. It simply ignores the possibility that money can be held offline and transacted online. Specifically, it is possible that a consumer would hold digital currency tokens in a non-custodial wallet and transfer them to a merchant with online connectivity, in much the same way that a consumer might rely upon the online connectivity of a POS device offered by the merchant.
There is a clear case for using digital currency with non-custodial wallets as a substitute for electronic card payments or network platform transactions at the point of sale. What remains is the question of whether there exists a need to support payments between two parties such that neither of which is connected to a network. There is no evidence that consumers who cannot withdraw cash or merchants who refuse to accept cash are using alternatives to cash that do not make use of a network connection instead, and there is no evidence of a significant demand for one, either. As a means of facilitating a transaction between two parties in situations in which neither party has a networked device, cash is remarkably effective. Although it is possible to imagine narrowly construed use cases for new capabilities, such use cases do not justify system requirements.
It is worth asking whether governments or central banks are harbouring an intention to do away with cash entirely, or whether the presumption of such an intention is shaping the view of what CBDC should be. For example, some authorities have argued that a larger share of domestic transactions should be brought inside the scope of the formal economy [15, 16]. Legitimate or not, that goal is not inconsistent with prevailing mechanisms to roll out digital transactions in general and does not require CBDC. Moreover, few monetary authorities or government regulators have called for the abolition of cash, whereas many, including those of Switzerland, the United Kingdom, and the United States, have pledged to preserve access to cash for the foreseeable future. Notably, Sweden has rolled back its drive toward creating a cashless society to support cash infrastructure in a near-cashless economy, at great cost [17]. So the idea that CBDC is mostly a justification for phasing out cash, as has been speculated by some, is not corroborated by facts.
3 Certified hardware as a de facto custodian
Design approaches based on certified hardware have featured prominently in the recent debate concerning the deployment of large-scale digital currency systems in general and systems to support CBDC in particular. Certified hardware is a cornerstone of the so-called “trusted computing” paradigm, wherein a user carries a device designed to operate in a manner that respects the will of a third party in preference to the will of the user. The justification for such approaches is usually given in terms of the perceived need to facilitate offline payments or to facilitate the recovery of lost assets. In this section, we examine the characteristics of trusted computing within the context of its application to digital currency systems and its potential impact on the power relationships between the users of devices, the manufacturers of devices, and other powerful actors such as corporations and the state.
Baressi and Zatti explicitly assume the use of trusted hardware to draw transactions into the security envelope of a trusted custodian [5], implicitly drawing the hardware manufacturer into the role of trusted third party. In their model, the wallet would be a means of transacting with money held in custodial accounts that would be reconciled ex post.
3.1 Fair exchange
If we define a fair exchange as a transaction mechanism that ensures that transacting parties either each receive or each do not receive what it expects (and do not receive “any additional information about the other’s item”) [19], then it is not possible to have a fair exchange between two parties without the involvement of a third party; this is a foundational truth [20]. With the exchange of physical objects, possession of the object itself is sufficient to demonstrate the veracity of the exchange that has taken place. But with digital objects the situation is more complicated. It is possible to have fair exchange without connection to the Internet, provided that there are third parties accessible locally, perhaps via a local network, as might be possible with ad hoc networks.
Proponents of approaches requiring certified hardware have proposed embedding into end-user devices a tamper-resistant “secure element” that can be used to systematically restrict what the device can do, disclose the immutable identity of the device to a third party, or furnish information to third parties about what the device had observed or done. The secure element would be impenetrable to the user, and the user would not be able to modify its behaviour without rendering the device inoperable. In effect, such an element effectively acts as the third party by proxy, by embedding the will of the third party into the operation of a device carried by a user.
3.2 The perils of trusted hardware
In the case of digital currency, the secure element would presumably be used as an on-board authority that would prevent double-spending by maintaining a record of transactions made by the device and disallowing users from transacting tokens that have already been signed over to others. This design model in particular, and trusted hardware in general, introduces a variety of problems. We consider some of those problems here:
  1. 1.
    Trusted computing introduces significant security risks particular to digital currency. The risks of relying upon the security of a hardware manufacturer are too great for a token-based CBDC context. In particular, a system design that relies upon trusted hardware components is fragile by design. If the trusted hardware fails to function as advertised or is compromised by an attacker, then users could potentially transact the same tokens an arbitrary number of times without bound. The system-level costs of such an attack would be huge, and the incentives would be tremendously enticing for attackers. We note that modern devices, such as debit cards, that rely upon tamper-proof hardware for the exchange of value generally interact with accounts. If their security is compromised, then the value available to attackers is limited by rules and operational procedures implemented by custodians. Without custodial accounts, such backup mechanisms are not possible, and so, a successful attack on the hardware root of trust would imply that it would be necessary for all users to replace their devices. Thus, it is reasonable to conclude that it is preferable to rely upon a protocol to achieve the required level of security, as can be done when one of the parties has access to the network, rather than to expect that trusted hardware modules would guarantee that all transactions would take place within the security envelope of a trusted custodian. To assume that all hardware wallets must take the latter approach because some circumstances might require it is tantamount to overspecification at the expense of users who would readily hold assets offline but transact them offline. It is worth noting that when Chaum, Grothoff, and Möser argue that “any economically producible device that stores tokens with monetary value in an individual’s possession, and enables offline transactions… will be the target of successful forgery attacks…” [34], they are referring to wallets that perform dual-offline transactions, in contrast to wallets that hold tokens offline and transact them online.
  2. 2.
    Secure elements are unaccountable to their users. Because trusted hardware is impenetrable to users, users cannot know with certainty what their devices are doing. The inability to analyse the behaviour of a device undermines the trust that users can and should place in the device. In essence, the device serves a second master that is not the user. Users cannot trust what they cannot verify, so they have every reason to be suspicious of unchecked promises on the part of hardware manufacturers or other privileged guarantors [21]. For example, secure elements can furnish information that can be used to identify the user or the user’s previous transactions. By contrast, with open hardware, skilled members of the community can conduct audits and provide reports to the public. Of course, users with the means to do so can personally conduct the audits themselves or hire contractors to conduct the audits for them. However, it is not necessary for every user to actually run an audit; evidence suggests that similar benefits are achieved so long as all users can perform audits if they choose to do so [22].
  3. 3.
    Certified hardware discourages innovation. Users who cannot create their own devices cannot innovate. Instituting a certification requirement for hardware serves to ensure that only those with the time and resources to pursue certification will be able to produce working devices. Regulations that increase the cost of entry are known to decrease entrepreneurship as well as innovation by existing firms [23]. Ordinary users and hobbyists will not be able to develop their own devices, creating a barrier to public understanding the of the core technology, encumbering public debate and consideration of alternative designs [24, 25].
  4. 4.
    Certified hardware distorts the market by privileging its manufacturers. The requirement to use trusted hardware effectively locks the hardware manufacturer in to a privileged position in the market, wherein prospective manufacturers require certification by some authority (either directly by the government or by a delegated authority such as a trade organisation), and concomitant lobbying and regulation can restrict free entry, increasing market concentration [26]. Furthermore, even if all manufacturers could have certification at zero marginal cost, the fact that fabrication facilities carry high fixed costs means that the market is likely to be highly concentrated, with manufacturers capturing the surplus [27].
  5. 5.
    A certified device is tantamount to a relationship with a custodian. Hardware manufacturers become gatekeepers at best, and globally trusted third parties in the general case. But unlike trusted third parties on the network, users bind themselves to long-term relationships with personal devices that contain trusted hardware; the relationships are not much different between account-holders and custodians, wherein account-holders are subject to satisfying the requirements of custodians. The cost of switching bank accounts is generally only a matter of time and effort on the part of account-holders, whereas the cost of switching devices generally includes per-device manufacturing and retailing cost plus taxes and costs associated with physical acquisition and disposal of the devices. In addition, while it is possible to implement both technical protocols and regulations that rely upon locally-trusted third parties, economies of scale mean that hardware manufacturers are likely to have global reach. It is reasonable to conclude that relationships between users and their devices featuring specific trusted hardware will be sticky indeed.
For these reasons, we are unconvinced that trusted hardware solutions are appropriate as requirements for the design of tools that allow users to possess and control their own assets. At the same time, we would not suggest that trusted hardware should be forbidden. In some fraction of cases in which physical cash is not an option and network connections are not available, trusted hardware might be an alternative kind of third party for transacting parties to consider, although this possibility is not sufficient justification to institute a requirement. In all cases, both sender and recipient must willingly choose to trust the third party to consummate the transaction, and given the final three points in the last above, it is difficult to imagine that a user preference for a particular device manufacturer would be real. Consider, for example, that existing users of secure hardware in the form of electronic payment cards generally have little say in which payment network their issuing banks choose.
3.3 Standards and lobbying
It is plain to see why a hardware manufacturer would lobby in favour of regulations and standards that declare “offline transactions” to be an essential requirement for CBDC. But this argument is unsound: Cash is an excellent offline payment method, and to our knowledge, few if any central banks have suggested that the purpose of CBDC is to replace cash outright. The argument implicitly equivocates between a concept of “offline” that denotes transactions wherein a user is not required to carry a device with network connectivity and a concept of “offline” that involves two counterparties physically meeting in a location without third parties or network infrastructure of any kind nearby. The former scenario is the only one in which users commonly use certified hardware today. The idea that cash could suddenly become unavailable for use in the latter scenario is an overly simplistic justification for hardware manufacturers to tout their products.
We note that secure elements can be used for a variety of purposes, including the prevention of exfiltration of private data from a device. However, requiring the use of secure elements for the purpose of creating attestations to parties other than the user of the device (i.e. “remote attestations”) that a device is operating within the security envelope of a trusted third party, whether a hardware manufacturer, service provider, or authority, implies a specific model for securing transactions that is not necessary for single-offline payments. For the reasons described above, we argue that single-offline payments would be both more robust and more secure if counterparties can trust network service providers, and the various methods network service providers can use to demonstrate the veracity of their claims, rather than having no choice but to rely upon trusted hardware physically held by the other counterparty. And so, for single-offline payments, trusted hardware adds nothing, but imposes significant costs, and therefore it must not be considered a requirement. Dual-offline transactions can certainly make use of remote attestation if both parties agree, but such transactions should not be seen as a general case for digital currency transactions, particularly given pervasive connectivity and the security benefits of using protocol-based security rather than introducing a requirement for counterparties to trust each other’s devices. For this reason, the development of standards that recommend or require hardware wallets to use a secure element for remote attestation are certainly problematic.
Standards that recommend hardware wallets to use a secure element to prevent the exfiltration of private data might be appropriate in the case of transactions involving smartphones, because of their complexity and large attack surface. However, it is inappropriate to assume that all or most digital wallets will be smartphones, not least because of their intrinsic security and privacy risks, despite being the “benchmark in user-accessible features”, as Barresi and Zatti argue [5].
4 Digital currency means tokens, not accounts
Barresi and Zatti assume that wallets would represent end-user money as account balances [5]. However, for ordinary consumers and non-financial businesses to hold digital currency directly, there must be some way for digital currency to be held outside accounts. Accounts generally involve a relationship between a beneficial owner (the account holder and a fiduciary, typically a regulated financial institution. A salient property of accounts is that access to assets held within accounts is determined by the identity of the account holder [28]. In contrast, tokens are assets that can be held outside of accounts. Tokens can be bearer instruments, meaning that they can be used in a way that does not index the identity of their owners. Payments can be held offline and transacted online, so the user experience would be similar to an EMV card payment except in which the payer presents digital assets rather than instructions to debit an account.
Compared to assets held within accounts, assets that are held directly by their owners have a different set of security requirements for their owners. For example, owners of assets held within accounts can rely upon various identification mechanisms, such as PIN numbers and two-factor authentication, to protect their assets if their cards or mobile devices are stolen, and asset custodians can also provide credit to account holders, such that if assets are spent unlawfully, the account holders can ask the asset custodians cover the value of their losses. In contrast, the bearer is responsible for bearer instruments. If cash is lost or stolen, owners generally have little recourse beyond law enforcement mechanisms to recover their money, and such mechanisms often fail to recover the value. The potential for loss or theft exists not only physically but also electronically. Devices can malfunction, and devices can be compromised. The prevailing model for handling devices that no longer work or have been compromised is to wipe their contents completely or to replace them wholesale. When such devices are used to access accounts, such replacement is not much of a problem, although bearer instruments inexorably expose their owners to new risks that require different solutions.
Some central banks, including the ECB [14], the US Federal Reserve [29], and the Bank of England [31], have presented arguments that holding limits for wallets are necessary for financial stability and to prevent bank runs. Such arguments are nonsense. Limiting CBDC withdrawals, rather than aggregate holdings, would be no less effective in preventing bank runs, and the approach of limiting bank withdrawals has been used successfully in the past, notably in Greece in 2015 [30]. Importantly, limiting CBDC withdrawals can be done without requiring wallets to implement account-like control mechanisms that undermine the will of their users. Barresi and Zatti suggest that wallets could perform “behavioural analytics” on their users [5]; little could be more disrespectful of a user’s autonomy.
Although a third party is necessary for fair exchange (and indeed trusted hardware or closed-source software blobs can be this third party de facto), Pagnia and Gärtner have shown that the role of the third party can be limited, even zero-knowledge: tokens can be transferred without the third party knowing what is being transferred or having a custodial relationship with either party [20].
The idea that digitalisation implies the use of accounts is not true in general. Digital tokens can be real indeed, with ownership demonstrable by transferring it: using a single-use private key to transfer the asset to another single-use private key. The requirement of an account in the form of a “wallet provider” is not true for cryptocurrencies in general, nor does a private key constitute a proof of identity when it is created to be used exactly once. It is true that many Bitcoin users rely on service providers to manage their holdings. However, to argue from this that self-custody of tokens will never take off is pure speculation, and it ignores the fact that most Bitcoin users are speculators with no expectation that they would use Bitcoin for ordinary retail purposes. Many people carry around an amount of cash that is useful for retail purchases but small enough that they could lose it without too much regret, despite the fact that losing it is forever.
Digital assets can certainly be held by users without accounts of any kind, for example, as a payer can in Chaum’s model. Accounts for recipients might be desirable, but they are not technically necessary. The involvement of a third-party to prevent double-spending does not imply that the third party must have a role as custodian or account provider, or knowledge of the asset being transferred, the circumstances of the transfer, or the identity of either party.
Tokens are also important for privacy. Systems that make use of transactions involving balances, even if the transactions themselves are privacy-preserving [36, 37], intrinsically require end-users to retain data produced during a transaction that can be used to prove that the transaction was carried out according to the rules. This means that transactions are not independent of one another. The property of transaction independence is important because retaining data related to past transactions introduces the risk that the user could be de-anonymised via poor operational security or blackmailed to reveal information about those transactions [38].
Finally, the idea that digital currency intrinsically requires accounts is not true; central banks can certainly issue tokens that exist outside of accounts, just as they do with cash, and a token like this would be a liability on a central bank’s balance sheet, not unlike cash or central bank reserves.
5 Separating issuance from transaction processing
Even if transactions are done with network access to a mutually trusted third party, there is no reason to assume that the issuer must be involved in transaction processing in real-time. Chaum’s original proposal assumed such a dependency [33], but this is not a requirement in general. In particular, it is possible to register a (zero-knowledge) commitment to transfer an asset without communicating with the issuer, and for a transaction to have finality it is sufficient for both parties have faith in a mutually trusted third party to eventually post the commitment to a bulletin board recognised by the issuer [10]. Also, if tokens carry their own provenance information, then there is no need for the issuer to track tokens that it had seen, either [10].
It is certainly possible to transact money issued by a central bank in a decentralised system, and wallets can be expected to respect this possibility. Trusting a central bank (or other issuing authority, such as those that issue e-money) to issue money and maintain its integrity is certainly different from trusting it to process transactions. In particular, it is possible to use a decentralised system to transact centrally issued money. The use of a secure element to establish trust implicitly requires the transacting parties to trust a hardware manufacturer or other central authority to facilitate transactions. Such a role for a unified central authority might be incompatible with the concept of a decentralised system. Finally, requirements for hardware wallets used for digital currency should not be needlessly inconsistent with requirements for hardware wallets used for other digital assets.
It is also worth considering that all users cannot be expected to trust the same trust anchors equally. Although the fact that assets held in offline non-custodial wallets can be transacted online is sufficient to reject the requirement of a hardware root of trust for security, even the case of dual-offline transactions poses a problem for requiring the use of secure elements. Specifically, there is the matter of who the transacting parties choose to trust. We imagine that parties choosing to transact online could potentially choose from a variety of manufacturers of hardware roots of trust based upon their needs and security models, whereas forcing manufacturers of hardware wallets to embed hardware certified by a specific authority, with the expectation that this authority is universally and equally trusted by all users, is not appropriate. We believe that allowing users to retain a choice of mutually trusted hardware security solutions, such as removable eSIM devices, is appropriate, although it might be impractical for users to carry around as many hardware devices as there are commonly trusted third parties with their various counterparties. The fact that hardware devices impose a trust model based upon their specific manufacturer is yet another reason to reject the idea that such devices could function as a public utility.
6 Wallets as containers for bearer instruments
For a set of specific technical requirements for digital currency in general to be credible, it must address the requirements of its various stakeholders. One of the unfortunate characteristics of the current debate is the diversity of conflicting assumptions about the nature of payments and the relative importance of different kinds of payment scenarios. In reality, there are many different ways that people make payments today: with and without custodians, locally and at a distance, with and without electronic devices, with and without money issued by a central bank. Money is transferred to pay invoices, to purchase goods and services, to give gifts, to pay taxes, or to reallocate funds within an organisation or family. These different scenarios imply a plethora of different costs and risk profiles: Buying a cup of coffee is different from buying a house. We should accept that there will not be a single payment system or key mechanism to rule them all, although it might be possible to develop a common framework that can accommodate a variety of different payment systems.
The case for trusted hardware, as for other core mechanisms of proposed payment systems, involves the requirement for assessing the integrity of money in a fair exchange scenario wherein counterparties lack access to a mutually trusted third party. In principle, trusted hardware can act as a trusted third party to enforce system-level rules on user devices directly, by implementing rules that prevent assets from being spent more than once. However, considering the security risks associated with compromised devices, it is better to enforce such rules over the network instead, if it is possible to do so. This can be done, for example, by invoking a trusted third party to validate all transactions, as proposed by Chaum and others [33, 34]. However, there is reason to believe the costs, risks, and public accountability concerns associated with having a central operator dictate that history unchecked are too great for the CBDC use case [35], and the most important costs and risks would be borne not only by the users but also by the central operator. A modern alternative involves using a distributed ledger system to put the money itself on a blockchain, the way that UTXO designs (e.g. Bitcoin [32], Monero [39]) and state-transition designs (e.g. Ethereum [40]) do. Such approaches are unwieldy because they require the ledger to track all of the tokens individually, and transaction finality requires global consensus. In practice, most cryptocurrency users rely upon custodial accounts, negating a foundational benefit of digital currency.
However, distributed ledgers and blockchains can be used to support and enable payment systems in a variety of ways, including by providing non-repudiable evidence that a transaction had taken place, even if an individual service provider might equivocate [41]. The digital currency design proposed by Goodell et al [10] and further elaborated by Friolo et al [38] addresses the scalability problems associated with modern cryptocurrency systems while strongly preventing double spending without trusted hardware. In particular, it avoids requiring the issuer to participate in every transaction of its tokens, it avoids requiring system operators to manage a list of tokens that have been created or spent, and it avoids requiring users to have access to a global network at the time of the transaction. Any third party chosen by the spender and accessible to the recipient at the time of the transaction can suffice.
Proposed solutions that rely upon stateful assets and oblivious transfers can work technically without distributed ledgers, but in the end, digital currency systems require a means to support an immutable, shared history mutually accepted among independent actors, which is the salient characteristic of a distributed ledger [42]. The immutability of a distributed ledger is emergent property of the independence of the participants. So, although using a distributed ledger to record and process tokens directly on the ledger might bring unacceptable costs in terms of scale, efficiency, and operational burden, we should not rule out using distributed ledgers for what they do best: externalising commitments so that parties responsible for maintaining the history do not equivocate [10]. These commitments can be aggregated in such a way that the ledger grows at a fixed rate determined in advance, and each transaction carries zero marginal cost to the ledger operators.
Trusted computing is not necessary to protect users from loss or theft, either. To protect their assets, users may, at their option, choose devices that offer security features such as encryption, passcodes, or local biometrics to protect against theft, and they may make copies of their tokens to mitigate the risk of accidental loss of their devices. Users can retain copies of their tokens, and they can share copies of the tokens with the police if the devices containing them are lost, stolen, or compromised. Users can also upload copies of their tokens to non-custodial service providers, such as cloud storage services.
A much better way to address the limited set of cases in which “fully offline” payments (that is, those where neither the payer nor the payee have access to any suitable third parties or network infrastructure) are necessary would be to support the use of cash itself where cash continues to work well, such as in remote locations and environments with low connectivity, while avoiding all of the aforementioned problems associated with trusted hardware. To address the reduction of cash in the “built” environments by providing digital currency, whilst encouraging the provision of facilities for converting between cash and diital currency at the boundary of “unbuilt” environments, would be an entirely legitimate approach. In particular, if one potential benefit of CBDC in an age of deteriorating cash infrastructure is the ability to electronically transport cash to remote population centres during crises [43], then there is no reason to require that central bank money distributed in this manner must remain digital. For example, it can be exchanged for physical tokens specific to that population centre, much as cash and bank money are exchanged for chips at a casino.
None of this is to say that we should forbid limited groups of users from trusting certified devices, if that is what they freely choose, to conduct transactions without network connectivity to a third party. Such an approach is entirely compatible with an architecture that expects that the third party would generally operate as a network service. However, at a system level, requiring all users to purchase and carry certified devices, even if some users intend only to hold money offline so that they can transact it online, introduces costs and risks that are entirely avoidable.
7 Conclusion
In the end, we must ask ourselves: What problem are we really trying to solve with digital currency? Manifestly, the problem is the decline of cash, and the concomitant challenges of maintaining cash infrastructure when the variable revenues of cash infrastructure fall below the fixed costs. But the decline of cash really only exists in environments with high network connectivity, wherein dual-offline payments are not necessary. For environments with low network connectivity, the decline of cash has not occurred, and cash continues to be the best solution for such environments.
Many central banks, including the Bank of England, have affirmed a commitment to cash, while underscoring the risks associated with the decline of cash use in favour of online payment schemes such as card payment networks. Really, the use cases that digital currency payments should be targeting are cases for which at least one counterparty is already online. To argue for digital currencies with the premise that support for dual-offline payments are its principal source of added value mainly supports the view that such products are solutions in search of a problem.
The purpose of digital wallets (and indeed, digital currency in general) is not primarily about enabling transactions in which both parties are offline. The purpose of digital wallets is to allow users to hold assets directly, without a custodian. Put differently, this is a way for users to have private property that they can exchange. By facilitating such payments, digital currency can protect the interests of consumers in all cases, including fully networked ones. If the counterparty has network connectivity (for example, an electronic point of sale device), then transactions involving digital wallets can be consummated without special hardware. Even if we accept that fair exchange is impossible without the participation of a trusted third party, and that the trusted third party can participate via the network or via trusted hardware, it is important to underscore that a non-networked digital wallet can conduct a fair exchange with a networked counterparty, without the aid of special security mechanisms built into the wallet.
It is also a mistake to assume the primary motivation for CBDC is to allow a government to phase out cash, or that custodial wallets are sufficient for all scenarios except the narrow case in which neither the sender nor the receiver have a communication line to a third party. Part of the reason that cash is successful is its role as a bearer instrument. Credit systems actually predated cash, and cash emerged as a partial substitute [44]. What the world does not yet have is an institutionally supported version of cash for the digital economy, and a digital currency aspiring to fill this role will require a non-custodial wallet architecture.
Custodial wallets require network access as a rule, because consumers must be able to use their devices to access their custodial accounts. However, digital wallets do not imply a design choice between network access and certified hardware in the general case. In particular, single-offline payments can make use of non-custodial wallets without certified hardware to hold money offline and transact it online, leveraging the network connectivity provided by POS devices presented to customers by merchants. To deny the salience of this use of wallets is to impede the development of a technology that may one day become crucial for the digital economy. Ultimately, the digital wallets that support economic transactions should be seen not as access devices or proxies offered by custodians, but as tools freely chosen, and even built, by end-users to store and use digital money autonomously, on their own terms.
Acknowledgements
The author thanks Professor Tomaso Aste for his continued support for our work on digital currencies and digital payment systems. The author also thanks his many colleagues in the international standards development community for a series of vigorous arguments that were instrumental to reaching the crux of this most contentious debate. The author also acknowledge the Future of Money Initiative at University College London, the Systemic Risk Centre at the London School of Economics, EPSRC, and the PETRAS Research Centre EP/S035362/1 for the FIRE Project.